Performance is back to normal and everything has been stable for the past hour. If anyone is still seeing issues, please put in a ticket to the support center so we can investigate and review.
Nov 18, 05:16 EST
The network configs have been updated and the hypervisor is back into full service. We are currently monitoring performance to make sure everything looks good.
Nov 18, 04:30 EST
We have swapped the faulty hardware out and re-configured all the necessary configs to bring this hypervisor back online. We are still performing the maintenance and do expect a couple brief network disturbances but all VMs will stay online. Thank you for your patience.
Nov 18, 04:04 EST
We are continuing to work on this issue and will have more updates to come.
Nov 18, 03:37 EST
We have identified the issue and are performing emergency maintenance work to fix the problem.
Nov 18, 02:48 EST
We are currently investigating an issue affecting one of our hypervisors.
Nov 18, 02:32 EST